# JEL classification: C31, J24

### v1

@@ 3   The Econometric Model and Identification Strategy
If effort were directly observable, we could estimate the peer effect parameter from Equation (3) following Kelejian and Prucha (1998). However, as we do not observe effort, the equation cannot be estimated directly. Instead, we derive from this equation an estimable equation based on GPA, which is influenced by effort. We show that this equation is econometrically different from the equation that we obtain if we proxy effort using GPA in Equation (2). We also present an identification strategy to identify the peer effect parameter $\lambda$ in the effort equation.
@@ 3.1 Reduced-Form Equation for GPA
From Equation (1), we express effort as a function of GPA and replace this expression in Equation (3). This yields a reduced-form equation for GPA that does not directly depend on effort (see Appendix A.2). The equation is given by
$y_{s,i}=\kappa_{s,i}+\lambda\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\tilde{\beta}}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\tilde{\gamma}}+(\boldsymbol{\omega}_{s,i}-\lambda\mathbf{g}_{s,i})\boldsymbol{\eta}_{s}+\delta^{2}\varepsilon_{s,i},$ (4)
where $\kappa_{s,i}=\delta^{2}c_{s}+(1-\lambda\mathbf{g}_{s,i}\mathbf{1}_{n_{s}})\alpha_{s}$, $\boldsymbol{\tilde{\beta}}=\delta^{2}\boldsymbol{\beta}+\boldsymbol{\theta}$, $\boldsymbol{\tilde{\gamma}}=\delta^{2}\boldsymbol{\gamma}-\lambda\boldsymbol{\theta}$, and $\boldsymbol{\omega}_{s,i}$ is a row-vector of dimension $n_{s}$ in which all the elements are zero, except the $i$-th element, which is one.
@@
$y_{s,i}=\bar{c}_{s}+\bar{\lambda}\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\bar{\beta}}+\delta\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\bar{\gamma}}+\bar{\varepsilon}_{s,i},$ (5)
@@
Let $\mathcal{V}^{NI}_{s}$ denote the subsample of students of school $s$ who have friends (non-isolated) and $\mathcal{V}^{I}_{s}$ denote the subsample of students of school $s$ who have no friends (isolated). As $\mathbf{G}_{s}$ is row-normalized, we have $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=1$ if $i\in\mathcal{V}^{NI}_{s}$ and $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=0$ otherwise. Thus, $\kappa_{s,i}=\kappa^{I}_{s}$ if $i\in\mathcal{V}^{I}_{s}$ and $\kappa_{s,i}=\kappa^{NI}_{s}$ if $i\in\mathcal{V}^{NI}_{s}$, where $\kappa^{I}_{s}=\delta^{2}c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=\delta^{2}c_{s}+(1-\lambda)\alpha_{s}$. If there are no isolated students in the network, $\kappa_{s,i}$ would be a simple school fixed effect, and our framework would be equivalent to the classical model. [^7] An isolated student is a student who has no friends. However, this student may be a friend of others. We later refer to a fully isolated student as a student who has no friends and who is not a friend of others. The difference between the standard model and our framework is that the standard model has a single intercept term per school. In Equation (4), $\kappa_{s,i}$ accounts for unobserved school-level heterogeneity depending on whether the student $i$ has friends or not. We would have $\kappa^{NI}_{s}=\kappa^{I}_{s}$ if and only if $\alpha_{s}=0$, $\lambda=0$, or if every student has friends. These conditions may not hold in many settings.
@@
[…]
@@ 3.2 Identification and Estimation
As the effort $e_{s,i}$ is not observed, not all the parameters of the structural model can be identified. Specifically, $\delta$ cannot be identified because it only appears in Equation (4) through a product with other parameters. Similarly, $\boldsymbol{\theta}$ and $\boldsymbol{\gamma}$ cannot be identified separately, as they only appear in the expression of $\boldsymbol{\tilde{\gamma}}$. Nevertheless, we can identify the composite parameters $\boldsymbol{\tilde{\beta}}=\delta^{2}\boldsymbol{\beta}+\boldsymbol{\theta}$ and $\boldsymbol{\tilde{\gamma}}=\delta^{2}\boldsymbol{\gamma}-\lambda\boldsymbol{\theta}$, which are parameters in the reduced-form equation (4) that capture the causal effects of students’ characteristics $\mathbf{x}$ and average friends’ characteristics $\bar{\mathbf{x}}$ on GPA. Consequently, it is not necessary to identify $\delta$ and $\boldsymbol{\theta}$ to estimate these causal effects. [^10] However, the non-identification of $\delta$ poses challenges for implementing certain counterfactual analyses, particularly when examining direct shocks on GPA and preference shocks on student effort. Equation (4) suggests that an increase in $c_{s}$ by $\Delta c_{s}$ (preference shock) leads to a change of $\delta^{2}\Delta c_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})^{-1}\mathbf{1}_{n_{s}}$ in $\mathbf{y}_{s}$. The impact of this increase on GPA depends on $\delta$.
@@
Let $n^{NI}_{s}$ denote the number of students in school $s$ with peers, and $n^{I}_{s}$ the number of students in the school $s$ without peers. Let also $\boldsymbol{\ell}^{NI}_{s}=\mathbf{G}_{s}\mathbf{1}_{n_{s}}$ and $\boldsymbol{\ell}^{I}_{s}=\mathbf{1}_{n_{s}}-\boldsymbol{\ell}^{NI}_{s}$. Equation (4) can be written in matrix form at the school level as
$\mathbf{y}_{s}=\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}+\lambda\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{X}_{s}\boldsymbol{\tilde{\beta}}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\tilde{\gamma}}+(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\delta^{2}\boldsymbol{\varepsilon}_{s}.$ (6)
@@
Note that the number of unknown parameters to be estimated in Equation (6) grows to infinity with the number of schools (there are $2S$ dummy variables). This issue is known as an incidental parameter problem and may lead to inconsistent estimators (Lancaster, 2000). A common approach to consistently estimate the model is to eliminate the fixed effects $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. To do so, we define $\mathbf{J}_{s}:=\mathbf{I}_{n_{s}}-\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}-\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}$ and impose by convention that $\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}=0$ if $n^{I}_{s}=0$, and that $\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}=0$ if $n^{NI}_{s}=0$. Note that $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=\hat{n}_{s}$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{I}_{s}=n^{I}_{s}$, $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{I}_{s}=0$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. Thus, $\mathbf{J}_{s}\boldsymbol{\ell}^{I}_{s}=\mathbf{J}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. One can eliminate the term $\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}$ by premultiplying each term of Equation (6) by the matrix $\mathbf{J}_{s}$. [^11] By premultiplying each term by $\mathbf{J}_{s}$, we consider Equation (6) in deviation to the average within the student group, that is, $\mathcal{V}^{I}_{s}$ or $\mathcal{V}^{I}_{s}$. This eliminates the parameters $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. This implies that
$\mathbf{J}_{s}\mathbf{y}_{s}=\lambda\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{J}_{s}\mathbf{X}_{s}\boldsymbol{\tilde{\beta}}+\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\tilde{\gamma}}+\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\delta^{2}\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}.$ (7)
@@
The random term $\boldsymbol{v}_{s}:=\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\delta^{2}\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$ is comprised of two error vectors $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. To consistently estimate $\boldsymbol{\psi}:=(\lambda,\leavevmode\nobreak\ \boldsymbol{\tilde{\beta}}^{\prime},\leavevmode\nobreak\ \boldsymbol{\tilde{\gamma}}^{\prime})^{\prime}$, we impose the following assumption.
@@
Assumption 3.1 implies that $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ are exogenous with respect to $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. This suggests that there is no omission of important variables in $\mathbf{X}_{s}$, which are captured by $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. We later relax this assumption in our approach to controlling for network endogeneity by allowing for $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ to depend on $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ through unobserved factors to the econometrician (see Section 5).
@@
Identification in peer effects models can be challenging, particularly due to the reflection problem (Manski, 1993). Bramoullé et al. (2009) address this problem and provide necessary and sufficient conditions for identification. Their main condition requires that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. However, their approach does not apply to our framework because their identification results assume that there are no isolated students (see their main assumption in their Section 2.1). [^12] Bramoullé et al. (2020) also discuss the case involving isolated students. They argue that the presence of isolated students can help identify the peer effect parameter (see their Section 2.1.1). Recall that the two intercepts of our model are $\kappa^{I}_{s}=\delta^{2}c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=\delta^{2}c_{s}+(1-\lambda)\alpha_{s}$. The fact that only one intercept depends on $\lambda$ implies that $\lambda$ can be identified using isolated students if $\kappa^{I}_{s}=\kappa^{NI}_{s}$. However, as we allow GPA shocks $\alpha_{s}$ to be different from preference shocks $c_{s}$ in terms of their impact on effort, this condition is unlikely to hold. Consequently, the presence of isolated students does not help identify $\lambda$ in our case.
Given that existing identification results do not directly apply to our model, we extend the analysis in Bramoullé et al. (2009). We derive easy-to-verify conditions to address the reflection problem when the network includes students without friends.
@@
(i) $\lambda\boldsymbol{\tilde{\beta}}+\boldsymbol{\tilde{\gamma}}\neq 0$; (ii) There are students in the network separated by a link of distance three.
@@
Condition (i) is equivalent to stating that GPA is influenced by at least one contextual variable. [^13] See Equation A.7, which quantifies the total effect of an increase in a contextual variable on GPA. With several characteristics in $\mathbf{X}_{s}$, this condition can be satisfied. Condition (ii) is slightly stronger than the assumption that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. It means that there are students who have connections to students that extend to three degrees of separation—friends of friends of friends—who are neither directly their friends nor friends of their friends (see an illustration in Figure 1 below). As we allow isolated students, Assumption 3.2 is not a necessary condition for identification, as is the case in Bramoullé et al. (2009); rather, it is a sufficient condition that holds in many cases.
Under Assumption 3.2, we show that $\boldsymbol{\psi}$ is identified. We estimate $\boldsymbol{\psi}$ using a standard GMM approach. It is well known that the regressor $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}$ is endogenous in Equation (7). However, it can be instrumented by $\mathbf{J}_{s}\mathbf{G}_{s}^{2}\mathbf{X}_{s}$, as suggested by Kelejian and Prucha (1998) and Bramoullé et al. (2009). [^14] To avoid a weak instrument issue, the pool of instruments can also be expanded to $\mathbf{J}_{s}[\mathbf{G}_{s}^{2}\mathbf{X}_{s},\leavevmode\nobreak\ \dots,\leavevmode\nobreak\ \mathbf{G}_{s}^{p}\mathbf{X}_{s}]$ for some integer $p>2$. Let $\boldsymbol{\hat{\psi}}$ be the GMM estimator of $\boldsymbol{\psi}$. We have the following result.
Proposition 3.1.
[…]
The consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}$ are directly derived from Kelejian and Prucha (1998). We later discuss how the asymptotic variance of $\boldsymbol{\hat{\psi}}$ can be estimated.
We show that the reflection problem arises within a network with links of distance three only when variations in friends’ average characteristics have no influence on students’ outcomes, which would directly contradict Condition (i) of Assumption 3.2.151515Houndetoungan (2022) applies a similar approach to the case of nonlinear models. Below, we provide an intuition behind our identification approach and a formal proof in Appendix A.3. Figure 1 presents an example of a network with a link of distance three: $i_{2}$ is at three nodes from $i_{1}$; they are a friend of a friend of a friend of $i_{1}$. The reflection problem arises when $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ is perfectly collinear with $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$. For a non-isolated student, this collinearity is equivalent to saying that there exist vector of parameters, $\dot{\boldsymbol{\beta}}$ and $\dot{\boldsymbol{\gamma}}$, such that $\mathbb{E}(\mathbf{g}_{s,i}\mathbf{y}_{s}-y^{NI}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})=(\mathbf{x}_{s,i}-\mathbf{\hat{x}}_{s})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{g}_{s,i}\mathbf{X}_{s}-\bar{\mathbf{x}}^{NI}_{s})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y^{NI}_{s}$, $\mathbf{x}^{NI}_{s}$, and $\bar{\mathbf{x}}^{NI}_{s}$ are respectively the averages of $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ in $\mathcal{V}^{NI}_{s}$. The variables $\mathbf{g}_{s,i}\mathbf{y}_{s}$, $\mathbf{x}_{s,i}$, and $\mathbf{g}_{s,i}\mathbf{X}_{s}$ are taken in deviation with respect to their average in $\mathcal{V}^{NI}_{s}$ because of the matrix $\mathbf{J}_{s}$ that multiplies the terms of Equation (7). If we take the previous equation in difference between students $i_{1}$ and $i_{3}$, we obtain
As $i_{3}$ is $i_{1}$’s only friend and $i_{4}$ is $i_{3}$’s only friend, we have $\mathbf{g}_{s,i_{1}}\mathbf{y}_{s}=y_{s,i_{3}}$, $\mathbf{g}_{s,i_{3}}\mathbf{y}_{s}=y_{s,i_{4}}$, $\mathbf{g}_{s,i_{1}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{3}}$, and $\mathbf{g}_{s,i_{3}}\mathbf{X}_{s}=\mathbf{x}_{s,i_{4}}$. Thus Equation (8) implies $y_{i_{3}}^{e}-y_{i_{4}}^{e}=(\mathbf{x}_{s,i_{1}}-\mathbf{x}_{s,i_{3}})^{\prime}\dot{\boldsymbol{\beta}}+(\mathbf{x}_{s,i_{3}}-\mathbf{x}_{s,i_{4}})^{\prime}\dot{\boldsymbol{\gamma}}$, where $y_{i}^{e}=\mathbb{E}(y_{s,i}|\mathbf{G}_{s},\mathbf{X}_{s})$. As $\mathbf{x}_{s,i_{2}}$ does not appear in the previous equation, we can say that $y_{i_{3}}^{e}-y_{i_{4}}^{e}$ is independent of $\mathbf{x}_{s,i_{2}}$, conditional on $\mathbf{x}_{s,i_{1}}$, $\mathbf{x}_{s,i_{3}}$ and $\mathbf{x}_{s,i_{4}}$. Put differently, an increase in $\mathbf{x}_{s,i_{2}}$(ceteris paribus) has no impact on $y_{i_{3}}^{e}-y_{i_{4}}^{e}$, which means that the increase has either no impact on $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ or the same impact on $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$. The first implication is not possible because $\mathbf{x}_{s,i_{2}}$ is the vector of contextual variables for $i_{4}$ and Condition (i) of Assumption 3.2 implies that GPA is influenced by at least one contextual variable. Moreover, $\mathbf{x}_{s,i_{2}}$ cannot influence $y_{i_{3}}^{e}$ and $y_{i_{4}}^{e}$ in the same way because $i_{2}$ is a direct friend of $i_{4}$ and not directly linked to $i_{3}$. This is in contradiction to equation (8), therefore, $\mathbb{E}(\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}|\mathbf{G}_{s},\mathbf{X}_{s})$ cannot be a linear combination of $\mathbf{J}_{s}\mathbf{X}_{s}$ and $\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}$.

### v2

@@ 3  The Econometric Model and Identification Strategy
If effort were directly observable, we could estimate the peer effect parameter from Equation (3) following kelejian1998generalized. However, as we do not observe effort, the equation cannot be estimated directly. Instead, we derive from this equation an estimable equation based on GPA, which is influenced by effort. We show that this equation is econometrically different from the equation that we obtain if we proxy effort using GPA in Equation (2). We also present an identification strategy to identify the peer effect parameter $\lambda$ in the effort equation.
@@ 3.1 Reduced-Form Equation for GPA
From Equation (1), we express effort as a function of GPA and replace this expression in Equation (3). This yields a reduced-form equation for GPA that does not directly depend on effort (see Appendix LABEL:append:GPA). The equation is given by
$y_{s,i}=\kappa_{s,i}+\lambda\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\eta_{s,i}-\lambda\mathbf{g}_{s,i}\boldsymbol{\eta}_{s}+\varepsilon_{s,i},$ (4)
where $\kappa_{s,i}=c_{s}+(1-\lambda\mathbf{g}_{s,i}\mathbf{1}_{n_{s}})\alpha_{s}$.
@@
$y_{s,i}=\bar{c}_{s}+\bar{\lambda}\mathbf{g}_{s,i}\mathbf{y}_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\bar{\beta}}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\bar{\gamma}}+\bar{\varepsilon}_{s,i},$ (5)
@@
Let $\mathcal{V}^{NI}_{s}$ denote the subsample of students of school $s$ who have friends (non-isolated) and $\mathcal{V}^{I}_{s}$ denote the subsample of students of school $s$ who have no friends (isolated). As $\mathbf{G}_{s}$ is row-normalized, we have $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=1$ if $i\in\mathcal{V}^{NI}_{s}$ and $\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}=0$ otherwise. Thus, $\kappa_{s,i}=\kappa^{I}_{s}$ if $i\in\mathcal{V}^{I}_{s}$ and $\kappa_{s,i}=\kappa^{NI}_{s}$ if $i\in\mathcal{V}^{NI}_{s}$, where $\kappa^{I}_{s}=c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$. If there are no isolated students in the network, $\kappa_{s,i}$ would be a simple school fixed effect, and our framework would be equivalent to the classical model. [^7] An isolated student is a student who has no friends. However, this student may be a friend of others. We later refer to a fully isolated student as a student who has no friends and who is not a friend of others. The difference between the standard model and our framework is that the standard model has a single intercept term per school. In Equation (4), $\kappa_{s,i}$ accounts for unobserved school-level heterogeneity depending on whether the student $i$ has friends or not. We would have $\kappa^{NI}_{s}=\kappa^{I}_{s}$ if and only if $\alpha_{s}=0$, $\lambda=0$, or if every student has friends. These conditions may not hold in many settings.
@@
The main distinction between the standard model and Equation (4) arises from an omitted-variable problem in the standard model. This omitted variable is $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$ and may introduce a discrepancy between the peer-effect estimates of the two models. Since this omitted variable depends on the network matrix $\mathbf{G}_{s}$, it can induce correlation between the peer regressor $\mathbf{G}_{s}\mathbf{y}_{s}$ and the structural error term. Consequently, network-based instruments constructed using $(\mathbf{G}_{s},\mathbf{X}_{s})$ (see BramoulleDjebbariFortin2009) may violate the exclusion restriction. The sign of the bias in the peer effect parameter depends on the sign of the covariance between the instrument and the omitted variable $\lambda\alpha_{s}\mathbf{g}_{s,i}\mathbf{1}_{n_{s}}$, as well as on the sign of the covariance between the instrument and the instrumented variable $\mathbf{G}_{s}\mathbf{y}_{s}$. Determining this sign is not straightforward because $\mathbf{X}_{s}$ includes many variables, each potentially leading to an instrument that influences the omitted variable and $\mathbf{G}_{s}\mathbf{y}_{s}$ differently.
@@ 3.2 Identification and Estimation
Let $n^{NI}_{s}$ and $n^{I}_{s}$ denote the number of students in school $s$ with and without peers, respectively. In addition, let $\boldsymbol{\ell}^{NI}_{s}=\mathbf{G}_{s}\mathbf{1}_{n_{s}}$ and $\boldsymbol{\ell}^{I}_{s}=\mathbf{1}_{n_{s}}-\boldsymbol{\ell}^{NI}_{s}$. Equation (4) can be written in matrix form at the school level as
$\mathbf{y}_{s}=\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}+\lambda\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\boldsymbol{\varepsilon}_{s}.$ (6)
@@
Note that the number of unknown parameters to be estimated in Equation (6) grows to infinity with the number of schools (there are $2S$ dummy variables). This issue is known as an incidental parameter problem and may lead to inconsistent estimators. A common approach to consistently estimate the model is to eliminate the fixed effects $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. To do so, we define $\mathbf{J}_{s}:=\mathbf{I}_{n_{s}}-\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}-\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}$ and impose by convention that $\dfrac{1}{n^{I}_{s}}\boldsymbol{\ell}^{I}_{s}\boldsymbol{\ell}^{I\prime}_{s}=0$ if $n^{I}_{s}=0$, and that $\dfrac{1}{n^{NI}_{s}}\boldsymbol{\ell}^{NI}_{s}\boldsymbol{\ell}^{NI\prime}_{s}=0$ if $n^{NI}_{s}=0$. Note that $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=n^{NI}_{s}$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{I}_{s}=n^{I}_{s}$, $\boldsymbol{\ell}^{NI\prime}_{s}\boldsymbol{\ell}^{I}_{s}=0$, $\boldsymbol{\ell}^{I\prime}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. Thus, $\mathbf{J}_{s}\boldsymbol{\ell}^{I}_{s}=\mathbf{J}_{s}\boldsymbol{\ell}^{NI}_{s}=0$. One can eliminate the term $\kappa^{I}_{s}\boldsymbol{\ell}^{I}_{s}+\kappa^{NI}_{s}\boldsymbol{\ell}^{NI}_{s}$ by premultiplying each term of Equation (6) by the matrix $\mathbf{J}_{s}$. [^9] By premultiplying each term by $\mathbf{J}_{s}$, we consider Equation (6) in deviation to the average within the student group, that is, $\mathcal{V}^{I}_{s}$ or $\mathcal{V}^{I}_{s}$. This eliminates the parameters $\kappa^{I}_{s}$ and $\kappa^{NI}_{s}$. This implies that
$\mathbf{J}_{s}\mathbf{y}_{s}=\lambda\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{y}_{s}+\mathbf{J}_{s}\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{J}_{s}\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}.$ (7)
@@
The random term $\boldsymbol{v}_{s}:=\mathbf{J}_{s}(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})\boldsymbol{\eta}_{s}+\mathbf{J}_{s}\boldsymbol{\varepsilon}_{s}$ is comprised of two error vectors $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. To consistently estimate $\boldsymbol{\psi}:=(\lambda,\penalty 10000\ \boldsymbol{\beta}^{\prime},\penalty 10000\ \boldsymbol{\gamma}^{\prime})^{\prime}$, we impose the following assumption.
@@
Assumption 3.1 implies that $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ are exogenous with respect to $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. This suggests that there is no omission of important variables in $\mathbf{X}_{s}$, which are captured by $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$. We later relax this assumption in our approach to controlling for network endogeneity by allowing for $\mathbf{X}_{s}$ and $\mathbf{G}_{s}$ to depend on $\boldsymbol{\eta}_{s}$ and $\boldsymbol{\varepsilon}_{s}$ through unobserved factors to the econometrician (see Section LABEL:sec:endo).
@@
Identification in peer effects models can be challenging, particularly due to the reflection problem (m1993). BramoulleDjebbariFortin2009 address this problem and provide necessary and sufficient conditions for identification. Their main condition requires that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. However, their approach does not apply to our framework because their identification results assume that there are no isolated students (see their Section 3.1).101010bramoulle2020peer also discuss the case involving isolated students. They argue that the presence of isolated students can help identify the peer effect parameter (see their Section 2.1.1). Recall that the two intercepts of our model are $\kappa^{I}_{s}=c_{s}+\alpha_{s}$ and $\kappa^{NI}_{s}=c_{s}+(1-\lambda)\alpha_{s}$. The fact that only one intercept depends on $\lambda$ implies that $\lambda$ can be identified using isolated students if $\kappa^{I}_{s}=\kappa^{NI}_{s}$. However, as we allow GPA shocks $\alpha_{s}$ to be different from preference shocks $c_{s}$ in terms of their impact on effort, this condition is unlikely to hold. Consequently, the presence of isolated students does not help identify $\lambda$ in our case.
Given that existing identification results do not directly apply to our model, we extend the analysis in BramoulleDjebbariFortin2009. We derive easy-to-verify conditions to address the reflection problem when the network includes students without friends.
@@
(i) $\lambda\boldsymbol{\beta}+\boldsymbol{\gamma}\neq 0$; (ii) There are students in the network separated by a link of distance three.
@@
Condition (i) is equivalent to stating that GPA is influenced by at least one contextual variable. [^11] See Equation LABEL:eq:append:Dy, which quantifies the total effect of an increase in a contextual variable on GPA. With several characteristics in $\mathbf{X}_{s}$, this condition can be satisfied. Condition (ii) is slightly stronger than the assumption that $\mathbf{I}_{n_{s}}$, $\mathbf{G}_{s}$, $\mathbf{G}_{s}^{2}$, and $\mathbf{G}_{s}^{3}$ are linearly independent. It means that there are students who have connections to students that extend to three degrees of separation—friends of friends of friends—who are neither directly their friends nor friends of their friends (see an illustration in Figure 1 below).
Under Assumption 3.2, we show that $\boldsymbol{\psi}$ is identified. Before formally presenting this result and its proof, we first provide an illustration. We consider a network with a link of distance three as depicted in Figure 1: $i_{2}$ is at three nodes from $i_{1}$; they are a friend of a friend of a friend of $i_{1}$.